# 2] organization=Observatoire des Sciences de l’Univers Nantes Atlantique (Osuna), addressline=CNRS UAR-3281, Nantes Université, country=France

### v2

@@ Abstract
We report observations of stripe-like features in Enceladus’ plumes captured simultaneously by Cassini’s VIMS-IR and ISS NAC instruments during flyby E17, with similar patterns seen in VIMS-IR data from flyby E13 and E19. These parallel stripes, inclined at approximately 16° to the ecliptic and 43° to Saturn’s ring plane, appear continuous across images when projected in the J2000 frame. A bright stripe, most visible at wavelengths around 5 µm, acts as the zeroth-order diffraction peak of a reflection grating with an estimated groove spacing of 0.12–2.60 mm, while adjacent stripes are attributed to higher-order diffraction peaks. We suggest that this light-dispersing phenomenon originates from an inclined periodic structure within Saturn’s E ring. This structure, constrained between Saturn’s G-ring and Rhea’s orbit, likely consists of fresh ice particles supplied by Enceladus’ plumes.
@@ 1 Introduction
The plumes and E ring provide a unique opportunity to investigate Enceladus’ internal processes and evolutionary history. In-situ analysis of E ring particles with Cassini’s Cosmic Dust Analyzer (CDA) revealed the presence of ice enriched in sodium salts (Postberg et al., 2009). When combined with gravity field measurements (Park et al., 2024), these findings support the existence of a global subsurface ocean beneath Enceladus’ icy crust, from which the plumes originate. As a result, the plumes and E ring particles provide direct samples of this subsurface ocean. Subsequent research detected complex organic molecules in the icy grains of the plumes (Postberg et al., 2018; Choblet et al., 2022). These findings point to a potentially habitable environment beyond Earth, which makes Enceladus a prime target for future missions aimed at searching for extraterrestrial life (Carr et al., 2013; Porco et al., 2017). This emphasis is further supported by ESA’s Voyage 2050 program, which prioritizes the exploration of icy moons, including Enceladus, as part of its long-term strategy for Solar System exploration (Favata et al., 2021). Awaiting future missions, Cassini’s extensive observations from the ISS and VIMS instruments continue to yield new scientific insights nearly 20 years after its initial deployment, e.g. Morello and Berg (2024) and Denny et al. (2024).
@@ 2.3.2 ISS Mosaic
The mosaic reveals that stripe features are visible only in the last eight images of the E17 approach, with the final four corresponding to those shown in Figure 1. Enceladus’ South Pole is present in all 16 images in the top center of each image window. The stripes appear near-continuous across images taken with the same filter. However, there is a slight misalignment where the last image is shifted by $DEC\approx 0.1^{\circ}$ compared to the second-to-last image. This misalignment could result from small pointing errors in the SPICE kernels or parallax. Nevertheless, the intensity variations remain consistent across the images, as seen in the order and intensity of the bright and dark regions in the IR1 images.
@@ 2.3.6 Cassini Observation Geometry
An overview of the observation geometry of Cassini during flyby E17 is shown in Figure 6. Starting from the edge-on view of Saturn’s ring plane, it can be seen that at the beginning of the approach, Cassini is at a distance of 8.5 $\mathrm{R_{S}}$ (Saturn Radii) and at an altitude of 2600 km above the ring plane. During its approach, Cassini moves closer to Saturn and the ring plane. The observation window of the bright band observed by six VIMS cubes in Figure 4 is shown in red. The bright band appears for a range of $7.26.9<\mathrm{R_{S}}<$ 7.2, before crossing Dione’s orbit. Shortly after this crossing, the bright area appears and remains visible till the end of the imaging approach. A top-down view is shown on the right side of Figure 6, revealing the observed sections of Saturn’s inner moon system. Note that the Sun’s elevation above the ring plane is not zero but 13.7 North from the center of the ring plane. The bright stripe is visible when the F-ring should appear in the background of the field of view. In addition, the G-ring and other inner rings (indicated by the grey area) can also be present in the background of images after the bright band observations.
Figure 6: Cassini’s E17 flyby geometry on 27/03/2012 from 09:53:14 to 14:49:33 UTC, shown from a Saturn-centered perspective with edge-on (left) and top-down (right) views of Saturn’s ring plane. Cassini’s position is marked by a diamond, and the moons’ positions are indicated by circles. The open symbol represents the start position, and the filled symbol represents the end position. The orbits of Enceladus, Tethys, Dione, and Rhea (E, T, D, R) are represented by dashed grey lines, while the F-ring and G-ring (F, G) are shown with solid grey lines. The solar direction is indicated by a dash-dotted arrow, and the ISS/VIMS viewing direction by black dashed arrows.
@@ 2.5 Stripe Observation Summary
A summary of the top-down geometry of the VIMS stripe observations relative to the Saturn system and aligned with Saturn’s solar illumination is presented in Figure 7. All observations were conducted when Cassini was near the ring plane and within the extent of the E ring (from Mimas’ orbit to Titan’s (Horányi et al., 2008)). Flybys E17 and E19 exhibit very similar viewing geometries, with Cassini positioned between the orbits of Tethys and Rhea, facing Saturn. In contrast, flyby E13 features a different geometry, with Cassini located between the G-ring and Enceladus’ orbit, this time looking away from Saturn.
@@
From the flyby mosaic shown in Figure 4, it became apparent that the stripes are fixed in RA and DEC space. Yet, Enceladus’s location is not; it traces its orbit with time, as seen from the Enceladus-centered VIMS cubes (each cube contains Enceladus in the same area of the viewing window). Therefore, the observed stripe phenomenon occurs in a system decoupled from the motion of Enceladus and can thus not occur due to a physical structure which co-rotates with Enceladus orbit. This means that the origin of the stripes likely lies in front or behind Enceladus along Cassini’s line of sight.
@@ 3.1.1 Stripe Spectral Dependencies within Single ISS and VIMS Images
Figure 8: Intenstiy curves of ISS NAC sequence 12 (top) and VIMS cube 1711553290 (bottom) along a sample line indicated by the red arrows. The ISS $I/F$ values are normalised by CLEAR filter images. The VIMS $I/F$ samples are shown for spectral bands with varying central wavelengths and a bandwidth of 50 nm.
@@
The VIMS intensity curves for different wavelengths are examined at the bottom of Figure 8. The line samples are obtained using linear interpolation of the pixel values along the line and then processed through a moving average smoothing algorithm to achieve smooth intensity curves despite the limited number of available pixels. Noise in the data is further reduced by sampling the average intensity of spectral bands with a bandwidth of 50 nm (4 VIMS channels) instead of a single channel. Starting from the shortest wavelength at 2.81 µm in deep blue, a peak at a sampling line distance of 9 pixels appears. With increasing wavelength, this peak shifts further to the right along the sample line and decreases in intensity until 2.88 µm (shown in grey). Increasing the wavelength further, the intensity increases again, and the peak continues to shift towards the right. Extending the analysis to the neighbouring VIMS channels results in the same periodic intensity fluctuations and the continuous shift in stripe spatial position with increasing wavelengths. Additionally, the periodic intensity fluctuations follow a sinusoidal pattern, where the global intensity decreases with increasing wavelength while the amplitude remains constant. The direction of this spatial shift is the same for the ISS and VIMS images.
@@ 3.1.2 Retrieval of Spectral Structures Throughout The Flyby
This methodology can also be extended to an entire flyby mosaic, as illustrated starting from the top right of Figure 9. Initially, the mosaic is rotated, and then the column averages are retrieved for each wavelength channel. These averages are then stacked to generate a spectral kymograph mosaic. However, the horizontal physical dimension now depicts an angular distance projected onto the celestial sphere. Step D demonstrates how an individual image cube is integrated within the kymograph mosaic. In practice, the kymograph mosaic is constructed by superimposing each individual cube kymograph onto the next in a manner analogous to the creation of the cube mosaics. The intensity values in each cube kymograph are normalized over the entire flyby, such that an $I/F$ value of 0 corresponds to the global minimum, and an $I/F$ value of 1 represents the global maximum throughout the flyby.
@@ 3.2 Characterisation of the Material at the Origin of the Stripes
(A) 1–2 µm & 2–2.6 µm Spectral Slopes: The bright band spectrum is similar to the E ring background for $\lambda<$ 2 µm aside from two smooth peaks centered around 1 and 1.6 microns. The plume is generally brighter in this spectral range with a similar spectral slope as the bright band and the E ring. In the spectral range of 2–2.6 µm, the bright band spectrum has a steeper slope than the plume and E ring spectra. Because the wavelength range of the observations is related to the particle sizes in the scene, the spectrum of the bright band suggests a particle size distribution with a higher percentage of larger particles compared to the plumes and E ring background (Hedman et al., 2009; Dhingra et al., 2016).
@@ 4 Interpretation & Discussion
A single bright line observed by VIMS-IR during flyby E17 lies at the center of these stripe observations. This bright band becomes visible at $\lambda>3$ µm and increases in contrast up until $\lambda=$ 5.11 µm. The width of this bright band is approximately 1 to 1.5 mrad, which is similar to the apparent width of the Sun as observed by Cassini. It was found that this feature acts as a reflection grating where the bright band itself is observed as reflected light, while the other stripes in the flyby are likely created from higher-order diffraction peaks. The estimated groove spacing of the grating should be 0.12 to 2.60 mm according to the diffraction effects seen in the bottom right plot of Figure 9.
The spectrum of the bright band’s material suggests that it consists primarily of crystalline water ice. The particle size distribution likely contains larger particles than the plume and E ring background or a lack of particles with a radius smaller than 2 µm. Additionally, a trace of solid phase CO2 was identified in the bright band’s material, which was not seen in the other spectra. These findings suggest that the material could be fresh and minimally affected by space-weathering.
Hypothesis: A millimetre-sized inclined ordered structure within the E ring. The stripes are produced by an organised structure (acting as a reflection grating) positioned between Cassini and Enceladus; as for the stripes to appear in front of Enceladus, they must be produced by particles located physically in front of the moon. For flybys E17 and E19, this structure lies specifically between the orbits of Enceladus and Rhea, and for E13, between the G-ring and Enceladus’ orbit. This range of radial distances from Saturn spans the E ring’s inner and outer regions. A visual representation of the periodic E ring structure is given in Figure 12.
Figure 12: Visualisation of the hypothesis for the observed stripe features, with a reflection grating located at the E ring. A zoomed-in view is shown in the bottom right corner.
The fact that the stripe features are visible only when observing Enceladus and occur both interior and exterior to its orbit suggests that Enceladus is likely the source of the stripe phenomenon. The bright band’s spectrum indicates the presence of crystalline water-ice and CO2 that resembles fresh plume material. Dhingra et al. (2017) and the analysis in subsection 3.2 show that the plumes consist mainly of crystalline water ice while Brown et al. (2006) and Combe et al. (2019) detected CO2 trapped in ice on the surface of Enceladus near the Tiger Stripes. However, the spectra of the plume and bright band’s material in Figure 11 remain different, leaving questions about the connection between the bright band’s material and the fresh plume material. If the bright band’s material is similar to fresh plume material, the plume spectrum would be expected to exhibit the CO2 absorption feature as well. However, the plume’s signal is noisier compared to the bright band at these wavelengths (as shown on the bottom left in Figure 11), which may obscure the CO2 absorption in the plume spectrum. From subsection 3.2, we noted that spectral differences could be attributed to particle size distributions favouring larger particles in the bright band compared to the E ring and the plume. If the bright band indeed consists of fresh plume ice, why are its particles larger? Three possible explanations arise: (1) the bright band ice is more shielded from space weathering than ice sampled directly from the plume, (2) crystal growth occurs within the E ring, or (3) a process selectively removes the smallest particles in the plumes, leaving behind only the larger ones.
We observed that the bright band is inclined approximately 43 to Saturn’s ring plane. Kempf et al. (2010); Hedman et al. (2011) found that E ring particles launched from Enceladus’ plumes can achieve orbital inclinations up to 1, nowhere near the required inclination of the stripes. Ice crystals may become preferentially oriented with the electric field generated by Saturn’s magnetosphere (Saunders and Rimmer, 1999; Jones et al., 2006), potentially creating the necessary inclination. This alignment can also lead to the formation of unusual halo displays (Moilanen and Gritsevich, 2022). However, the intricate interactions between the plumes of Enceladus and Saturn’s magnetosphere, along with the uncertainty surrounding the shapes of the plume ice crystals, complicate predictions regarding halo phenomena and crystal orientations. In addition, these ice particle halos require ice crystals that are significantly larger than the observed wavelength (Bi and Yang, 2014). However, most particles in the E ring are comparable in size to the observed wavelengths of the ISS and VIMS (Hedman et al., 2009; Ingersoll and Ewald, 2011), making diffraction the primary mechanism for light scattering in this context. This aligns with the diffraction behaviour observed in Figure 9. In addition to opaque ice being periodically arranged to form a diffraction grating, diffraction effects may also result from the shape of the ice crystals, similar to the behaviour observed in photonic crystals (Joannopoulos et al., 1997).
Mitchell et al. (2015) found sinusoidal structures within the E ring near Enceladus called ’tendrils’. These tendrils originate from the highest velocity particles ejected from the plumes and follow horseshoe orbits from Enceladus’ leading and trailing edge. These horseshoe orbits’ looping motion could potentially produce the necessary inclination and periodic structure for the stripe features. Simulations from Weiss et al. (2013) concluded that for particles smaller than $\sim$1 micron, the influence of Saturn’s magnetic field blurs out the tendrils while larger particles $\sim$3 microns create distinct tendrils. This preference for larger particles aligns with the spectra shown in Figure 11.
@@
Another much smaller source of particles in the E ring comes from the impacts of existing E ring particles on the inner mid-sized moons (Dobrovolskis et al., 2010). These impacts eject fresh surface material to the E ring. Kempf et al. (2008) simulated orbits from these impact ejecta and found low inclinations up to 1 similar to those originating from the plumes. Other material, such as that from Saturn’s inner rings, is also unlikely to produce high inclinations since it remains confined to Saturn’s ring plane. The close alignment of the stripes with the ecliptic could hint towards a source beyond Saturn’s Hill sphere, like interplanetary dust or cometary debris. However, such contributions are relatively small compared to material generated within Saturn’s system (Hsu et al., 2011).

### v3

@@ Abstract
We report observations of stripe-like features in Enceladus’ plumes captured simultaneously by Cassini’s VIMS-IR and ISS NAC instruments during flyby E17, with similar patterns seen in VIMS-IR data from flyby E13 and E19. These parallel stripes, inclined at approximately 16° to the ecliptic and 43° to Saturn’s ring plane, appear continuous across images when projected in the J2000 frame. A bright stripe, most visible at wavelengths around 5 µm, acts as the zeroth-order diffraction peak of a reflection grating with an estimated groove spacing of 0.12–2.60 mm, while adjacent stripes are attributed to higher-order diffraction peaks. We suggest that this light-dispersing phenomenon originates from an inclined periodic structure within Saturn’s E ring. This structure, constrained between Saturn’s G ring and Rhea’s orbit, likely consists of fresh ice particles supplied by Enceladus’ plumes.
@@ 1 Introduction
The plumes and E ring provide a unique opportunity to investigate Enceladus’ internal processes and evolutionary history. In-situ analysis of E ring particles with Cassini’s Cosmic Dust Analyzer (CDA) revealed the presence of ice enriched in sodium salts (Postberg et al., 2009). When combined with gravity field measurements (Park et al., 2024), these findings support the existence of a global subsurface ocean beneath Enceladus’ icy crust, from which the plumes originate. As a result, the plumes and E ring particles provide direct samples of this subsurface ocean. Subsequent research detected complex organic molecules in the icy grains of the plumes (Postberg et al., 2018; Choblet et al., 2022). These findings point to a potentially habitable environment beyond Earth, which makes Enceladus a prime target for future missions aimed at searching for extraterrestrial life (Carr et al., 2013; Porco et al., 2017). This emphasis is further supported by ESA’s Voyage 2050 program, which prioritises the exploration of icy moons, including Enceladus, as part of its long-term strategy for Solar System exploration (Favata et al., 2021). Awaiting future missions, Cassini’s extensive observations from the ISS and VIMS instruments continue to yield new scientific insights nearly 20 years after its initial deployment, e.g. Morello and Berg (2024) and Denny et al. (2024).
@@ 2.3.2 ISS Mosaic
The mosaic reveals that stripe features are visible only in the last eight images of the E17 approach, with the final four corresponding to those shown in Figure 1. Enceladus’ South Pole is present in all 16 images in the top centre of each image window. The stripes appear near-continuous across images taken with the same filter. However, there is a slight misalignment where the last image is shifted by $DEC\approx 0.1^{\circ}$ compared to the second-to-last image. This misalignment could result from small pointing errors in the SPICE kernels or parallax. Nevertheless, the intensity variations remain consistent across the images, as seen in the order and intensity of the bright and dark regions in the IR1 images.
@@ 2.3.6 Cassini Observation Geometry
An overview of the observation geometry of Cassini during flyby E17 is shown in Figure 6. Starting from the edge-on view of Saturn’s ring plane, it can be seen that at the beginning of the approach, Cassini is at a distance of 8.5 $\mathrm{R_{S}}$ (Saturn Radii) and at an altitude of 2600 km above the ring plane. During its approach, Cassini moves closer to Saturn and the ring plane. The observation window of the bright band observed by six VIMS cubes in Figure 4 is shown in red. The bright band appears for a range of $7.26.9<\mathrm{R_{S}}<$ 7.2, before crossing Dione’s orbit. Shortly after this crossing, the bright area appears and remains visible till the end of the imaging approach. A top-down view is shown on the right side of Figure 6, revealing the observed sections of Saturn’s inner moon system. Note that the Sun’s elevation above the ring plane is not zero but 13.7 North from the centre of the ring plane. The bright stripe is visible when the F ring should appear in the background of the field of view. In addition, the G ring and other inner rings (indicated by the grey area) can also be present in the background of images after the bright band observations.
Figure 6: Cassini’s E17 flyby geometry on 27/03/2012 from 09:53:14 to 14:49:33 UTC, shown from a Saturn-centred perspective with edge-on (left) and top-down (right) views of Saturn’s ring plane. Cassini’s position is marked by a diamond, and the moons’ positions are indicated by circles. The open symbol represents the start position, and the filled symbol represents the end position. The orbits of Enceladus, Tethys, Dione, and Rhea (E, T, D, R) are represented by dashed grey lines, while the F ring and G ring (F, G) are shown with solid grey lines. The solar direction is indicated by a dash-dotted arrow, and the ISS/VIMS viewing direction by black dashed arrows.
@@ 2.5 Stripe Observation Summary
A summary of the top-down geometry of the VIMS stripe observations relative to the Saturn system and aligned with Saturn’s solar illumination is presented in Figure 7. All observations were conducted when Cassini was near the ring plane and within the extent of the E ring (from Mimas’ orbit to Titan’s (Horányi et al., 2008)). Flybys E17 and E19 exhibit very similar viewing geometries, with Cassini positioned between the orbits of Tethys and Rhea, facing Saturn. In contrast, flyby E13 features a different geometry, with Cassini located between the G ring and Enceladus’ orbit, this time looking away from Saturn.
@@
From the flyby mosaic shown in Figure 4, it became apparent that the stripes are fixed in RA and DEC space. Yet, Enceladus’s location is not; it traces its orbit with time, as seen from the Enceladus-centred VIMS cubes (each cube contains Enceladus in the same area of the viewing window). Therefore, the observed stripe phenomenon occurs in a system decoupled from the motion of Enceladus and can thus not occur due to a physical structure which co-rotates with Enceladus orbit. This means that the origin of the stripes likely lies in front or behind Enceladus along Cassini’s line of sight.
@@ 3.1.1 Stripe Spectral Dependencies within Single ISS and VIMS Images
Figure 8: Intensity curves of ISS NAC sequence 12 (top) and VIMS cube 1711553290 (bottom) along a sample line indicated by the red arrows. The ISS $I/F$ values are normalised by CLEAR filter images. The VIMS $I/F$ samples are shown for spectral bands with varying central wavelengths and a bandwidth of 50 nm.
@@
The VIMS intensity curves for different wavelengths are examined at the bottom of Figure 8. The line samples are obtained using linear interpolation of the pixel values along the line and then processed through a moving average smoothing algorithm to achieve smooth intensity curves despite the limited number of available pixels. Noise in the data is further reduced by sampling the average intensity of spectral bands with a bandwidth of 50 nm (4 VIMS channels) instead of a single channel. Starting from the shortest wavelength at 2.81 µm in pale pink, a peak at a sampling line distance of 9 pixels appears. With increasing wavelength, this peak shifts further to the right along the sample line and decreases in intensity until 2.88 µm (shown in orange). Increasing the wavelength further, the intensity increases again, and the peak continues to shift towards the right. Extending the analysis to the neighbouring VIMS channels results in the same periodic intensity fluctuations and the continuous shift in stripe spatial position with increasing wavelengths. Additionally, the periodic intensity fluctuations follow a sinusoidal pattern, where the global intensity decreases with increasing wavelength while the amplitude remains constant. The direction of this spatial shift is the same for the ISS and VIMS images.
@@ 3.1.2 Retrieval of Spectral Structures Throughout The Flyby
This methodology can also be extended to an entire flyby mosaic, as illustrated starting from the top right of Figure 9. Initially, the mosaic is rotated, and then the column averages are retrieved for each wavelength channel. These averages are then stacked to generate a spectral kymograph mosaic. However, the horizontal physical dimension now depicts an angular distance projected onto the celestial sphere. Step D demonstrates how an individual image cube is integrated within the kymograph mosaic. In practice, the kymograph mosaic is constructed by superimposing each individual cube kymograph onto the next in a manner analogous to the creation of the cube mosaics. The intensity values in each cube kymograph are normalised over the entire flyby, such that an $I/F$ value of 0 corresponds to the global minimum, and an $I/F$ value of 1 represents the global maximum throughout the flyby.
@@ 3.2 Characterisation of the Material at the Origin of the Stripes
(A) 1–2 µm & 2–2.6 µm Spectral Slopes: The bright band spectrum is similar to the E ring background for $\lambda<$ 2 µm aside from two smooth peaks centred around 1 and 1.6 microns. The plume is generally brighter in this spectral range with a similar spectral slope as the bright band and the E ring. In the spectral range of 2–2.6 µm, the bright band spectrum has a steeper slope than the plume and E ring spectra. Because the wavelength range of the observations is related to the particle sizes in the scene, the spectrum of the bright band suggests a particle size distribution with a higher percentage of larger particles compared to the plumes and E ring background (Hedman et al., 2009; Dhingra et al., 2016).
@@ 4 Interpretation & Discussion
A single bright line observed by VIMS-IR during flyby E17 lies at the centre of these stripe observations. This bright band becomes visible at $\lambda>3$ µm and increases in contrast up until $\lambda=$ 5.11 µm. The width of this bright band is approximately 1 to 1.5 mrad, which is similar to the apparent width of the Sun as observed by Cassini. The 43 inclination of this feature to Saturn’s ring plane is difficult to reconcile with known physical structures in the E ring. Kempf et al. (2010); Hedman et al. (2012) found that E ring particles launched from Enceladus’ plumes can only reach orbital inclinations up to 1. Impacts of E ring particles on the inner mid-sized moons also contribute material to the E ring (Dobrovolskis et al., 2010), but simulations show similarly low inclinations (Kempf et al., 2008). Other material, such as that from Saturn’s inner rings, is also unlikely to produce high inclinations since it remains confined to Saturn’s ring plane. The close alignment of the stripes with the ecliptic could hint towards a source beyond Saturn’s Hill sphere, like interplanetary dust or cometary debris. However, such contributions are relatively small compared to material generated within Saturn’s system (Hsu et al., 2011).
Ice crystals may become preferentially oriented with the electric field generated by Saturn’s magnetosphere (Saunders and Rimmer, 1999; Jones et al., 2006), potentially creating the necessary inclination. This alignment can also lead to the formation of unusual halo displays (Moilanen and Gritsevich, 2022). However, the intricate interactions between the plumes of Enceladus and Saturn’s magnetosphere, along with the uncertainty surrounding the shapes of the plume ice crystals, complicate predictions regarding halo phenomena and crystal orientations. In addition, these ice particle halos require ice crystals that are significantly larger than the observed wavelength (Bi and Yang, 2014). However, most particles in the E ring are comparable in size to the observed wavelengths of the ISS and VIMS (Hedman et al., 2009; Ingersoll and Ewald, 2011), making diffraction the primary mechanism for light scattering in this context. This aligns with the diffraction behaviour observed in Figure 9. We found that the stripes display chromatic behaviour similar to a reflection grating where the bright band itself is observed as reflected light, while the other stripes in the flyby are likely created from higher-order diffraction peaks. The estimated groove spacing of the grating should be 0.12 to 2.60 mm.
The spectrum of the bright band’s material suggests that it consists primarily of crystalline water ice. The particle size distribution likely contains larger particles than the plume and E ring background or a lack of particles with a radius smaller than 2 µm. Additionally, a trace of solid phase CO2 was identified in the bright band’s material, which was not seen in the other spectra. These findings suggest that the material could be fresh and minimally affected by space-weathering. Dhingra et al. (2017) and the analysis in subsection 3.2 show that the plumes consist mainly of crystalline water ice while Brown et al. (2006) and Combe et al. (2019) detected CO2 trapped in ice on the surface of Enceladus near the Tiger Stripes. However, the spectra of the plume and bright band’s material in Figure 11 remain different, leaving questions about the connection between the bright band’s material and the fresh plume material. If the bright band’s material is similar to fresh plume material, the plume spectrum would be expected to exhibit the CO2 absorption feature as well. However, the plume’s signal is noisier compared to the bright band at these wavelengths (as shown on the bottom left in Figure 11), which may obscure the CO2 absorption in the plume spectrum. We noted that other spectral differences could be attributed to particle size distributions favouring larger particles in the bright band compared to the E ring and the plume. If the bright band indeed consists of fresh plume ice, why are its particles larger? Three possible explanations arise: (1) the bright band ice is more shielded from space weathering than ice sampled directly from the plume, (2) crystal growth occurs within the E ring, or (3) a process selectively removes the smallest particles in the plumes, leaving behind only the larger ones.
@@
Hypothesis: A millimetre-sized inclined ordered structure within the E ring. The stripes are produced by an organised structure (acting as a reflection grating) positioned between Cassini and Enceladus; as for the stripes to appear in front of Enceladus, they must be produced by particles located physically in front of the moon. For flybys E17 and E19, this structure lies specifically between the orbits of Enceladus and Rhea, and for E13, between the G ring and Enceladus’ orbit. This range of radial distances from Saturn spans the E ring’s inner and outer regions. The fact that the stripe features are visible only when observing Enceladus and occur both interior and exterior to its orbit suggests that Enceladus is likely the source of the stripe phenomenon. A visual representation of the periodic E ring structure is given in Figure 12. We identify two potential processes within the E ring that could give rise to the formation of the observed periodic structure.
Figure 12: Visualisation of the hypothesis for the observed stripe features with either density waves or ice aggregates located at the E ring. Periodic structures in these features act as a reflection grating with groove spacing $\lambda$ (we suggest mm-scale) as shown by the zoomed-in view in the bottom right corner.
1) Periodic dust density waves: E ring particles are organised into alternating high-density (more opaque) and low-density (more transparent) regions, due to external processes. Mitchell et al. (2015) found high-density sinusoidal structures within the E ring near Enceladus called ’tendrils’. These tendrils originate from the highest velocity particles ejected from the plumes and follow horseshoe orbits from Enceladus’ leading and trailing edge. These horseshoe orbits’ looping motion could potentially produce the necessary inclination and periodic structure for the stripe features. Simulations from Weiss et al. (2013) concluded that for particles smaller than $\sim$1 micron, the influence of Saturn’s magnetic field blurs out the tendrils while larger particles $\sim$3 microns create distinct tendrils. This preference for larger particles aligns with the spectra shown in Figure 11. Hedman and Young (2021) also demonstrated that similar high-density wave structures could result from rebound wakes, generated by gravitational perturbations on existing E ring particles due to Enceladus’ gravity.
Pickett et al. (2015) detected electrostatic solitary waves (ESWs) in data from Cassini’s Radio and Plasma Wave Science Wideband Receiver (WBR) attributed to plasma activity within Saturn’s magnetosphere. These ESWs were observed inside 10 $\mathrm{R_{S}}$, particularly near Enceladus between 4 and 6 $\mathrm{R_{S}}$—the same region where stripe-like features were observed during flyby E17. ESWs can drive dust acoustic waves (DAWs) in the E ring’s dusty plasma, producing alternating high- and low-density regions that may appear as the observed chromatic stripes. DAWs have been modelled near Enceladus by Singh et al. (2022), with predicted spatial spacings down to 11 meters. Laboratory experiments presented by Merlino (2009), though not specific to Saturn’s environment, demonstrate that dust acoustic waves can form at millimetre wavelengths in the presence of micron-sized dust particles—consistent with our observational constraints. This supports the physical plausibility of ultrasonic-scale DAWs forming within Saturn’s magnetosphere. Cassini’s WBR may have missed these fine-scale structures due to its limited sampling frequency, which, based on our rough estimate, resolves features no smaller than a few centimetres.
2) Periodic ice grain structure: beyond ice grains being periodically arranged, diffraction effects may also result from the shape of the ice crystals themselves, similar to the behaviour observed in photonic crystals (Joannopoulos et al., 1997). In this case, icy aggregates in the E ring could create the required millimetre-scale periodicity. For example, Gao et al. (2016) proposed that aggregates could form within Enceladus’ plumes. It is also possible that the high-density waves mentioned earlier might promote particle aggregation. Yet, the optical behaviour of such photonic crystal-like aggregates is complex and highly sensitive to their structure, making it difficult to predict their signatures or directly relate them to the observed diffraction features.